Calculate (a) the value of the resistor Rx, such that the total power dissipated in the circuit is 2.5KW, and (b) the current flowing in each of the four resistors . R1=15ohms, R2=10ohms, R3=38ohms and R4=Rx.
1
Expert's answer
2020-04-21T10:57:57-0400

As per the given question,

Here value of V and the diagram was not given, so I assumed that the given diagram is like given below,




"R_1=15\Omega, R_2=10\Omega, R_3=38\Omega R_4=R_x."

a) P=VI

"I=\dfrac{2.5\times 1000}{250}=10A"

"\Rightarrow 250=10(\dfrac{15\times 10}{15+10})+10\times\dfrac{38\times R_x}{38+R_x}"

"\Rightarrow 25=\dfrac{150}{25}+\dfrac{38R_x}{38+R_x}"

"\Rightarrow 25=6+\dfrac{38R_x}{38+R_x}"

"\Rightarrow \dfrac{38R_x}{38+R_x}=19"

"\Rightarrow 38R_x=38\times 19+R_x\times 19"

"\Rightarrow 19R_x=38\times 19"

"\Rightarrow R_x=38\Omega"